solve the following equations for the unknown:
6.2(x-15.1)=9.3(-x-3.2)and x/3-7/8=2/5(x+15/4)
any help would be greatly appreciated, thanks.

Answer by stanbon(75887)

(x-15.1)=9.3(-x-3.2)
----
x - 15.1 = -9.3x + 29.76
10.3x = 44.86
-----
x = 4.3553
=====================
and x/3-7/8=2/5(x+15/4)
Multiply thru by 120 to get:
40x - 105 = 48(x+(15/4))
---
40x - 105 = 48x + 180
8x = -285
x = 144.375
==================
Cheers,
Stan H.
===============